# Gaussian Process Tilted Nonparametric Density Estimation using Fisher Divergence Score Matching

### v1

@@ III-C Variational Predictive Distribution
Because the Fisher divergence consists of terms linear and quadratic in $\theta$, as shown in Section III-A, the solution to $q(\theta)$ is the following multivariate Gaussian,
$\qquad q(\theta)$ $=\mathcal{N}(\theta|\widehat{\mu},\widehat{\Sigma})$ (30)
$\widehat{\Sigma}$ $=\textstyle\left(\lambda I+\frac{1}{\eta\gamma^{2}}ZZ^{\top}\odot\Phi^{\prime}\right)^{-1}$
@@
We use the definitions of $\Phi^{\prime}$, $\psi$ and $\psi^{\prime}$ in Algorithm 2. We also observe a similar parallel between ridge regression and Bayesian linear regression from the perspective of $\theta$ in the relationship between $\widehat{\mu}$ and the solution to Algorithm 2, which has a MAP interpretation in this context.
Let $X=\{x_{1},\dots,x_{N}\}$. The predictive distribution is
[…]
The integral of this variational predictive distribution is intractable because $\theta$ appears in the intractable denominator of $q_{\theta}(x)$. To this end, we introduce the following MGF approximation of the normalizing constant of $q_{\theta}(x)$,
$\mathbb{E}_{x\sim\mathcal{N}(x|\mu,\Sigma)}\big{[}\exp\{\theta^{\top}\phi(x)\}\big{]}\approx\exp\{\theta^{\top}\mu_{\phi}+\frac{1}{2}\theta^{\top}\Sigma_{\phi}\theta\},$
@@
Using similar derivations as discussed in Section III-B for the noise conditional model, we can show that
$\mu_{\phi}(s)=$ $~{}\textstyle\exp\{-\frac{1}{2\gamma^{2}}z_{s}^{\top}\Sigma z_{s}\}\phi_{s}(\mu)$ (32)
$\Sigma_{\phi}=$ $~{}\textstyle\frac{1}{2}(\Delta^{-}+\Delta^{+})\odot\phi(\mu)\phi(\mu)^{\top}$ (33)
$+$ $~{}\textstyle\frac{1}{2}(\Delta^{-}-\Delta^{+})\odot\phi^{\prime}(\mu)\phi^{\prime}(\mu)^{\top}-\mu_{\phi}\mu_{\phi}^{\top}$
$\Delta_{ss^{\prime}}^{+}\equiv$ $~{}\textstyle\exp\{-\frac{1}{2\gamma^{2}}(z_{s}+z_{s^{\prime}})^{\top}\Sigma(z_{s}+z_{s^{\prime}}\}$
$\Delta_{ss^{\prime}}^{-}\equiv$ $~{}\textstyle\exp\{-\frac{1}{2\gamma^{2}}(z_{s}-z_{s^{\prime}})^{\top}\Sigma(z_{s}-z_{s^{\prime}})\}.$
@@
[…]
which can be solved up to a proportionality constant. After integrating, the approximate predictive distribution is proportional to
$q(x|X)\propto\mathcal{N}(x|\mu,\Sigma)\mathrm{e}^{\frac{1}{2}\phi(x)^{\top}\mathbf{M}^{-1}\phi(x)-\phi(x)^{\top}\mathbf{M}^{-1}\mathbf{m}},$
@@
We summarize our Fisher variational predictive distribution (FVPD) in Algorithm 4.
Model LCD CA Bikeshare FICO HIGGS MIMIC2 Credit Adult Churn
(5D, 10K) (8D, 14K) (12D, 11K) (36D, 7K) (28D, 50K) (17D, 16K) (30D, 182K) (14D, 21K) (19D, 5K)
FVPD -0.751$\pm$0.003 -6.62$\pm$0.02 -1.97$\pm$0.00 -3.59$\pm$0.17 -3.99$\pm$0.02 -2.77$\pm$0.01 -4.60$\pm$0.03 -2.09$\pm$0.01 -2.57$\pm$0.01
NCFD -0.752$\pm$0.003 -6.75$\pm$0.00 -1.96$\pm$0.01 -4.17$\pm$0.05 -4.00$\pm$0.02 -2.77$\pm$0.01 -4.61$\pm$0.02 -2.10$\pm$0.01 -2.56$\pm$0.01
FD -0.764$\pm$0.003 -7.26$\pm$0.03 -2.00$\pm$0.01 -3.63$\pm$0.19 -4.01$\pm$0.02 -2.94$\pm$0.01 -4.61$\pm$0.03 -2.10$\pm$0.01 -2.67$\pm$0.01
MAP -0.752$\pm$0.003 -6.43$\pm$0.01 -1.97$\pm$0.01 -3.98$\pm$0.06 -4.02$\pm$0.02 -2.79$\pm$0.01 -4.64$\pm$0.03 -2.10$\pm$0.01 -2.56$\pm$0.01
TABLE II: Average test log likelihood for the inference approaches described in Algorithms 1-4 for some common benchmark data sets. Dimensionality and training data size are indicated below the data name. Noise conditional and variational methods improve Fisher divergence learning, with our proposed variational predictive marginal likelihood often better than the noise conditional option. MAP performs well using this metric due to its similar objective function.
Algorithm 4 Fisher variational predictive TGP distribution
@@
5:Using all data, calculate $\Phi^{\prime}$, $\psi^{\prime}$, $\psi$ as in Algorithm 2.
6:Calculate $\widehat{\mu}$, $\widehat{\Sigma}$, $\mu_{\phi}$, $\Sigma_{\phi}$ (Eqs. 30, 32 & 33)
7:Construct the approximate predictive distribution
@@
(a) NCFD
(b) FVPD
(c) FD
(d) MAP
(e) KDE (RFF)
(f) KDE ($\infty$)
Figure 1: Samples (orange) shown over data (blue) generated from (a) the noise conditional Fisher divergence (Algorithm 3), (b) the Fisher variational predictive distribution (Algorithm 4), (c) Fisher divergence (Algorithm 2), (d) MAP inference (Algorithm 1), (e,f) kernel density estimation (see text for discussion). We set $S=1000$ for (a)-(d) and $S=100K$ for (e), $\lambda=10$ for (a,b,d) and $\lambda=100$ for (c). All methods use the same kernel width $\gamma$.
Figure 2: Noise conditional density contours on CA House data for decreasing $\sigma$. For better visualization, Figure 1 shows samples from the final RHS density contour (NCFD), along with samples from the other learned methods.
III-C1 Discussion
We can observe some properties of Equation 34 to connect it with Algorithm 2 and help guide setting of $\eta$. First we note that, since $\gamma^{2}$ can be very small, $\mathbf{M}^{-1}\mathbf{m}\approx-\widehat{\mu}$ and $\phi(x)^{\top}\mathbf{M}^{-1}\phi(x)\approx 0$ when $\eta=1$. Effectively, the predictions of Algorithm 2 and Algorithm 4 are the same since $\widehat{\mu}=\theta_{\mathrm{FD}}$. In this scenario, we empirically observed nearly identical results as Algorithm 2 in terms of low-density inaccuracy, since a huge $1/\gamma^{2}$ scaling removes uncertainty in $\theta$. To eliminate this effect, the tempering parameter $\eta$ can act as a regularizer to balance the Fisher and KL divergence terms of the ELBO. We found that setting $\eta=1/\gamma^{2}$ can remove the undesirable impacts of $\gamma$ in Equation 30.
We also observe that the predictive distribution in Equation 34 can only be solved up to a normalizing constant, since the integral over $x$ remains intractable, which requires additional considerations similar to other unnormalized score models. One discretization option here is to sample $x_{i}\sim_{iid}\mathcal{N}(x|\mu,\Sigma)$ many times and approximate $p(x_{i}|X)\propto\exp\{\frac{1}{2}\phi(x_{i})^{\top}\mathbf{M}^{-1}\phi(x_{i})-\phi(x_{i})^{\top}\mathbf{M}^{-1}\mathbf{m}\}$ on these samples. For generative models where one dimension of $x$ is predicted conditioned on values for the others, a 1D grid approximation of Equation (34) can be made over the free dimension.
@@ IV Experiments
We evaluate the four algorithms in this paper on nine low dimensional benchmark data sets and three other data sets. These include LCD, CA House, Bikeshare, FICO, HIGGS, MIMIC2, Credit, Adult and Churn from the UCI repository, along with a projected MNIST data set and two simulated 2D data sets. The dimensionality and training size of the benchmark data are indicated in Table II, with several thousand additional data held out for testing. For all experiments, we set $S=1000$, $\gamma=\frac{1}{10}\sqrt{\mathrm{tr}(\mathbb{V}(x))}$, and $\lambda=10$, except for the vanilla Fisher divergence, where $\lambda=100$. We set $\mu$ and $\Sigma$ in the base normal distribution to be their empirical values from the data. For Fisher divergence we optimized these values after learning $\theta_{\mathrm{FD}}$ without reiterating. For the noise conditional model we set $\sigma_{\mathrm{max}}=\sqrt{\mathrm{tr}(\mathbb{V}(x))/D}.$
The four inference algorithms for the GP-tiled density are MAP, Fisher Divergence (FD), Noise Conditional Fisher Divergence (NCFD), and Fisher Variational Predictive Distribution (FVPD). Considering its similarity with Kernel Density Estimation, we also show a qualitative result comparing with $p(x|X)\propto\sum_{i=1}^{N}k(x,x_{i})$ (KDE-$\infty$) and its approximation $p(x|X)\propto\phi(x)^{\top}\big{[}\sum_{i=1}^{N}\phi(x_{i})\big{]}$ (KDE-RFF), which is very fast but inaccurate when $S$ is small.
We first show qualitative results on dimensions 7 and 8 from the CA House data set due to it’s unique real world pattern. In Figure 1 we show the original data with 10K samples generated from each learned density super imposed. (The RHS of Figure 2 better visualizes the raw data.) In general, we see that the GP-tilted model is able to learn complex densities with a single vector $\theta\in\mathbb{R}^{S}$. MAP and FD underperform NCFD and FVPD in opposite directions. Compared with the standard NCFD model, our proposed FVPD is able to learn sharper densities by modeling “noise” (uncertainty) in the posterior of $\theta$ rather than in the noise-added data itself. KDE-RFF essentially failed for smaller $S$ due to artifacts, which are still visible when $S$=100,000. The exact KDE also does well, but is based directly on the raw data. In Figure 2 we show the noise conditional density contours for (NCFD) as a function of decreasing $\sigma$.
(i) NCFD
(j) FVPD
(k) FD
(l) MAP
@@
Figure 3: Contour plots of GP-tilted density learned by each algorithm on three data sets. Data shown in light blue.
@@
In Figure 3 we show contour plots for the four learning approaches of our GP-tilted density using three other data sets: MNIST-60K projected onto its first two principal components, and XO and Smile synthetic data. For NCFD, we show the final data-level density not integrated over $\sigma$. Noticeable in FD algorithm are a tendency to learn high density values outside of the data region. This issue due to the lack of data to learn meaningful gradients is addressed as intended by the other Fisher approaches, although we observe that NCFD has issues with XO. We also note that MAP also performs well on these problems.
We also evaluate our algorithms quantitatively in Table II on 9 public benchmark data sets. We use the expected log (predictive) likelihood learned by the corresponding algorithm. Since all distributions can only be learned up to a proportionality constant, we approximate the log normalizer with $100K$ samples to allow for direct comparison. We note that KDE (not shown) performed worse generally, with KDE-RFF having significant issues with approximation artifacts. We observe that MAP generally performs very well, since it is based on the log likelihood and evaluated on test data very similar to training data. Still, it is often beaten by Fisher divergence. We also observe that the basic Fisher divergence is often significantly worse than our uncertainty-informed versions. Of these, uncertainty captured by the model’s posterior distribution (FVPD) can frequently improve upon uncertainty captured by noise added to the data and modeled by the likelihood-based score function (NCFD).

### v2

@@ III-C Fisher Variational Predictive Distribution
Because the Fisher divergence consists of terms linear and quadratic in $\theta$, as shown in Section III-A, the solution to $q(\theta)$ is the following multivariate Gaussian distribution,
[…]
$\widehat{\Sigma}$ $=\big(\lambda I+\tfrac{1}{\gamma^{2}\eta}ZZ^{\top}\odot\Phi^{\prime}\big)^{-1}$
@@
The definitions of $\Phi^{\prime}$, $\psi$ and $\psi^{\prime}$ are taken from Algorithm 2. We observe that the equivalence between $\widehat{\mu}$ above and the Fisher divergence solution given in Algorithm 2 parallels that between ridge regression and Bayesian linear regression. In effect, the variational Fisher distribution $q(\theta)$ adds Gaussian uncertainty to the $\theta_{\mathrm{FD}}$ point estimate.
Using the posterior approximation $q(\theta)$, we can approximate the predictive distribution. Let data $X=\{x_{1},\dots,x_{N}\}$. The variational predictive distribution is
$q(x|X)=\int q_{\theta}(x)q(\theta)\,d\theta$
and we recall that $q_{\theta}(x)\propto\exp\{{\theta^{\top}\phi(x)\}}\mathcal{N}(x|\mu,\Sigma)$. The integral is intractable because $\theta$ appears in the also-intractable denominator of $q_{\theta}(x)$. We introduce the following MGF approximation of the normalizing constant of $q_{\theta}(x)$,
$\mathbb{E}_{x\sim\mathcal{N}(x|\mu,\Sigma)}\big[\exp\{\theta^{\top}\phi(x)\}\big]\approx\exp\{\theta^{\top}\mu_{\phi}+\tfrac{1}{2}\theta^{\top}\Sigma_{\phi}\theta\},$
@@
Using a similar derivation as in Section III-B for the noise conditional model, we can show that
$\mu_{\phi}(s)=$ $~\exp\{-\tfrac{1}{2\gamma^{2}}z_{s}^{\top}\Sigma z_{s}\}\phi_{s}(\mu)$ (32)
$\Sigma_{\phi}=$ $~\tfrac{1}{2}(\Delta^{-}+\Delta^{+})\odot\phi(\mu)\phi(\mu)^{\top}$ (33)
$+$ $~\tfrac{1}{2}(\Delta^{-}-\Delta^{+})\odot\phi^{\prime}(\mu)\phi^{\prime}(\mu)^{\top}-\mu_{\phi}\mu_{\phi}^{\top}$
$\Delta_{ss^{\prime}}^{+}:=$ $~\exp\{-\tfrac{1}{2\gamma^{2}}(z_{s}+z_{s^{\prime}})^{\top}\Sigma(z_{s}+z_{s^{\prime}})\}$
$\Delta_{ss^{\prime}}^{-}:=$ $~\exp\{-\tfrac{1}{2\gamma^{2}}(z_{s}-z_{s^{\prime}})^{\top}\Sigma(z_{s}-z_{s^{\prime}})\}.$
@@
$q(x|X)\approx\mathcal{N}(x|\mu,\Sigma)\int\mathrm{e}^{\theta^{\top}(\phi(x)-\mu_{\phi})-\frac{1}{2}\theta^{\top}\Sigma_{\phi}\theta}\mathcal{N}(\theta|\widehat{\mu},\widehat{\Sigma})\,d\theta.$
The integral can be solved and the approximate variational predictive distribution is
$q(x|X)\approx\mathcal{N}(x|\mu,\Sigma)\mathrm{e}^{\frac{1}{2}\phi(x)^{\top}\mathbf{M}^{-1}\phi(x)-\phi(x)^{\top}\mathbf{M}^{-1}\mathbf{m}},$
@@
We summarize this Fisher variational predictive distribution (FVPD) in Algorithm 4.
Algorithm 4 Fisher variational predictive distribution for TGP
@@
2:Run Algorithm 2 steps 1 through 4.
3:Calculate $\widehat{\mu}$, $\widehat{\Sigma}$, $\mu_{\phi}$, $\Sigma_{\phi}$ (Eqs. 30, 32 & 33)
4:Construct the approximate predictive distribution
@@
Algorithm Bikeshare MIMIC2 MAGIC HIGGS Credit Power GasMod GasMix HARTH Twitter
MAP 2.25|\,|\,3.60 1.78|\,|\,2.78 7.05|\,|\,14.1 0.67|\,|\,1.21 2.12|\,|\,4.62 11.7|\,|\,23.1 7.17|\,|\,11.3 7.80|\,|\,13.7 10.1|\,|\,21.4 6.00|\,|\,12.0
FD 2.25|\,|\,3.82 1.78|\,|\,2.84 5.45|\,|\,12.1 0.60|\,|\,1.03 2.10|\,|\,4.56 11.3|\,|\,23.0 6.79|\,|\,11.1 8.23|\,|\,14.0 10.1|\,|\,21.2 6.15|\,|\,12.4
NCFD 2.29|\,|\,3.81 1.79|\,|\,2.88 5.44|\,|\,11.9 0.66|\,|\,1.21 2.05|\,|\,4.42 11.0|\,|\,21.4 7.05|\,|\,11.9 8.49|\,|\,13.9 10.1|\,|\,21.9 5.72|\,|\,13.2
FVPD 2.28|\,|\,3.70 1.70|\,|\,2.76 5.34|\,|\,11.9 0.65|\,|\,1.19 2.01|\,|\,4.46 11.4|\,|\,23.0 6.91|\,|\,10.8 8.43|\,|\,13.5 10.1|\,|\,21.1 5.87|\,|\,12.3
MAP 2.78|\,|\,4.74 2.17|\,|\,3.49 7.44|\,|\,16.1 0.73|\,|\,1.38 2.32|\,|\,5.06 12.7|\,|\,26.0 7.71|\,|\,12.9 8.39|\,|\,15.2 10.3|\,|\,21.8 6.38|\,|\,12.6
FD 2.73|\,|\,4.67 1.98|\,|\,3.13 6.04|\,|\,13.8 0.67|\,|\,1.24 2.29|\,|\,4.93 12.2|\,|\,25.4 7.33|\,|\,12.8 8.70|\,|\,16.1 10.2|\,|\,21.7 6.77|\,|\,13.3
NCFD 2.70|\,|\,4.57 2.01|\,|\,3.23 5.94|\,|\,13.3 0.74|\,|\,1.43 2.24|\,|\,4.83 11.9|\,|\,24.6 7.43|\,|\,13.0 8.85|\,|\,16.1 10.3|\,|\,21.7 6.44|\,|\,13.9
FVPD 2.62|\,|\,4.47 1.96|\,|\,3.12 5.88|\,|\,13.5 0.72|\,|\,1.37 2.23|\,|\,4.85 12.0|\,|\,25.0 7.32|\,|\,12.2 8.64|\,|\,15.2 10.2|\,|\,21.4 6.40|\,|\,12.8
TABLE II: Kolmogorov-Smirnov test scores and Wasserstein distances (KS|\,|\,WD) between the empirical distributions and the marginal along random one dimensional subspaces. All values are shown on scale $\times 10^{-2}$. The top shows the median and the bottom the mean of the KS/WD values over uniform projection distribution using a 500-sample Monte Carlo approximation.
We can observe some properties of (34) to connect it with the base FD Algorithm 2 and help guide setting the parameter $\eta$ that appears in $\widehat{\mu}$ and $\widehat{\Sigma}$. First, since the squared kernel width $\gamma^{2}$ may be very small, $\mathbf{M}^{-1}\mathbf{m}\approx-\widehat{\mu}$ and $\phi(x)^{\top}\mathbf{M}^{-1}\phi(x)\approx 0$ when $\eta=1$. The predictions of Algorithm 2 and Algorithm 4 become nearly the same since $\widehat{\mu}=\theta_{\mathrm{FD}}$ and the covariance $\widehat{\Sigma}$ is small. In this scenario, we empirically observed almost identical results as Algorithm 2 since a huge $1/\gamma^{2}$ scaling removes uncertainty in $\theta$. To eliminate this effect, the tempering parameter $\eta$ can act as a regularizer to balance the Fisher and KL divergence terms of the ELBO. We found that setting $\eta\propto 1/\gamma^{2}$ can remove the undesirable impact of $\gamma$ in (30).
@@ IV Experiments
We conduct an empirical evaluation to demonstrate our proposed methods on lower dimensional density estimation problems. We test the our four inference algorithms for the GP-tiled density: MAP, Fisher Divergence (FD), Noise Conditional Fisher Divergence (NCFD), and Fisher Variational Predictive Distribution (FVPD).
Datasets
We evaluate the four algorithms in this paper on ten data sets from the UCI repository and four other illustrative data sets. These include Bikeshare, MIMIC2, MAGIC, HIGGS, Credit, Power, GasMod, GasMix, HARTH and Twitter. Statistics for these data sets are given in Table III. We also illustrate using 2D-projected MNIST, two dimensions of CA Housing (from UCI), and two simulated 2D data sets.
Parameter settings
We set $S=1000$ and $\lambda=0.1$. We also follow “Scott’s Rule” to set $\gamma=N^{-1/(d+4)}\sqrt{\mathrm{tr}(\mathbb{V}(x))/d}$ [scott2015multivariate]. We set $\mu$ and $\Sigma$ in the base normal distribution to be their empirical values from the data. For the noise conditional model we set $\sigma_{\mathrm{max}}=\frac{1}{d}\sqrt{\mathrm{tr}(\mathbb{V}(x))}$ and $H=10$. For FVPD, we set $\eta=\frac{1}{\gamma^{2}}$. All values were obtained through rough parameter selection, but could be further optimized. While we discuss TGP in the context of a spherical kernel, we sample $z\sim\mathcal{N}(0,\Sigma_{z})$, where $\Sigma_{z}=d\cdot\Sigma/\mathrm{tr}(\Sigma)$ to incorporate data correlation in the kernel function, with slight performance improvement.
Quantitative Result
We show a quantitative evaluation on 10 public benchmark data sets in Table II. As a performance measure, we use the Kolmogorov-Smirnov (KS) test and the Wasserstein distance (WD) to quantify similarity between the learned and empirical distributions. Let $\mathcal{P}_{N}$ be the empirical CDF of data and $Q$ be the CDF of a learned distribution on the same data, both in $\mathbb{R}$. KS and WD measure discrepancy between these CDFs in one dimension as
$\mathrm{KS}(\mathcal{P}_{N},Q)$ $=$ $\textstyle\sup_{x}|\mathcal{P}_{N}(x)-Q(x)|,$
$\mathrm{WD}(\mathcal{P}_{N},Q)$ $=$ $\textstyle\int_{\mathbb{R}}|\mathcal{P}_{N}(x)-Q(x)|\,dx.$
Dataset N D MAP FD NCFD FVPD
Bikeshare 11,122 12 5.7m 0.18s 3.6s 0.18s
MIMIC2 15,684 17 5.7m 0.31s 9.2s 0.31s
MAGIC 19,020 10 5.7m 0.28s 5.1s 0.28s
HIGGS 50,000 28 5.7m 0.56s 7.3s 0.56s
Credit 182,276 30 5.7m 2.3s 30s 2.3s
Power 2,049,280 7 6m 22s 4.5m 22s
GasMod 3,843,160 18 6.4m 43s 9.8m 43s
GasMix 4,208,261 16 6.5m 45s 10m 45s
HARTH 6,461,328 6 7m 68s 13m 68s
Twitter 14,262,515 3 8m 2.5m 33m 2.5m
TABLE III: Data statistics and total running times.
(a) MAP (Algorithm 1)
(b) FD (Algorithm 2)
(c) NCFD (Algorithm 3)
(d) FVPD (Algorithm 4)
(e) KDE (RFF, $S=100K$)
(f) KDE ($S=\infty$)
(g) Data
Figure 2: Samples (orange) shown over data (blue) generated for (a-d) Algorithms 1-4, and (e,f) kernel density estimation (see text for discussion). We set $S=1000$ for (a)-(d) and $S=100K$ for (e), $\lambda=10$ for (a-d). All methods use the same kernel width $\gamma$.
Figure 3: Algorithm 3 example. Noise conditional density contours on CA House data for decreasing $\sigma$. For better visualization, Figure 2 shows samples from the final RHS density contour (NCFD), along with samples from the other learned methods.
Figure 4: Example CDFs of projections of MAGIC data used to calculate KS and WD. Data (blue) vs FVPD (red). Each plot corresponds to a different random $\bm{\mathrm{v}}$.
(k) MAP (Algorithm 1)
(l) FD (Algorithm 2)
(m) NCFD (Algorithm 3)
(n) FVPD (Algorithm 4)
(o) Data
@@
Figure 5: Contour plots of GP-tilted density learned by each algorithm on three data sets.
Since the data is $d$-dimensional, we compute these measures by first generating a direction $\bm{\mathrm{v}}$ uniformly distributed on the unit sphere in $\mathbb{R}^{d}$ and then calculate the marginal $\mathrm{KS}_{\bm{\mathrm{v}}}$, $\mathrm{WD}_{\bm{\mathrm{v}}}$ along the direction $\bm{\mathrm{v}}$ at $10^{4}$ grid points $\alpha\bm{\mathrm{v}}$ defined by scalars $\alpha$. We calculate the marginal distribution $\mathcal{P}_{N}$ by projecting the data on $\bm{\mathrm{v}}$, and calculate the marginal $Q$ along thin slices $\mathcal{B}_{\alpha\mathrm{v}}$ perpendicular to $\mathrm{v}$ around grid scalings $\alpha$ by approximating the unnormalized marginal $\int\bm{1}(x\in\mathcal{B}_{\alpha\mathrm{v}})f_{\mathrm{alg}}(x)\mathcal{N}(x|\mu,\Sigma)dx$ using 250K samples from the Gaussian base. Here, $f_{\mathrm{alg}}(x)$ is the exponential term in either (5) or (34), depending on the algorithm used. Sampling creates a discrete distribution proportional to $f_{\mathrm{alg}}(\hat{x}_{s})\delta_{\hat{x}_{s}}$ at sample point $\hat{x}_{s}$. The weighted marginal CDF can be found by projecting the $\hat{x}_{s}$ on $\bm{\mathrm{v}}$ and weighting by $f_{\mathrm{alg}}(\hat{x}_{s})$. We use 500 samples of $\bm{\mathrm{v}}$ and center the data for these experiments.
In Table II, we see that MAP often performs worse than at least one FD-based learning method. Also, noise-conditional and variational methods often improve upon the base FD algorithm. While NCFD and FVPD are competitive with each other, FVPD is roughly $H$ times faster than NCFD ($H=10$ here; see Table III). In Figure 4 we show 10 example CDFs of random projections for MAGIC data using FVPD.
In Table III we show total run times for each algorithm. MAP is calculated over 10K iterations and is primarily a function of MC sample size rather than data size or dimension. We sample 100K points once and use the same samples for each iteration, significantly speeding up the MAP algorithm. FD and FVPD are both faster than NCFD roughly by a factor of parameter $H$, and since FVPD uses statistics calculated for FD with almost no extra overhead, they have the same run time. We can therefore see that FVPD has a significant time advantage as the data size grows.
Qualitative Results
We also illustrate TGP estimation on two dimensional data. First, we focus on dimensions 7 and 8 from the CA House data set due to it’s unique pattern in a real world data set. In Figure 2 we show the original data with 10K samples generated from each learned density superimposed. In general, we see that the GP-tilted model is able to learn complex densities with a single vector $\theta\in\mathbb{R}^{S}$ where $S=1000$. MAP and FD qualitatively underperform NCFD and FVPD in opposite directions; FD overfits, while MAP underfits. Compared with the standard NCFD model, our proposed FVPD is able to learn sharper densities. We conjecture it may achieve this because it models “noise” (uncertainty) in the posterior of $\theta$ rather than in the noise-added data itself. As baselines, we show classic kernel density estimation (KDE) on this data, for both exact and RFF approximation, according to
$p(x|X)~\propto~\frac{1}{N}\sum\nolimits_{i=1}^{N}k(x,x_{i})~\approx~\phi(x)^{\top}\Big[\frac{1}{N}\sum\nolimits_{i=1}^{N}\phi(x_{i})\Big].$
KDE-RFF performed poorly for smaller $S$ due to artifacts, which are still visible when $S=100,000$. The exact KDE also does well, but is based directly on the raw data and is computationally less appealing than a parameterized representation. In Figure 3 we also show the noise conditional density contours for (NCFD) as a function of decreasing $\sigma$, where we can see the multiscale advantage of that learning approach not available to FVPD.
@@
In Figure 5 we show contour plots for our TGP learning algorithms using three other toy data sets: MNIST-60K projected onto its first two principal components, and XO and Smile synthetic data. For NCFD, we show the final data-level density for $\sigma=0$. Noticeable in the FD algorithm is a tendency to learn high density values outside of the data region. This issue is addressed as intended by the other Fisher approaches, although we observe that NCFD has issues with XO. In these low dimensional toy data sets, we see that TGP is able to learn fairly complex densities without requiring a similarly complex density model. We also refer to Figure 4 for qualitative examples of how the learned densities in higher dimensions match the data by comparing their CDFs over 1D marginals.